how do you get the solution to 4(y - 2) = 2(y + 4) + 2
Answer by Gogonati(855)


4(y-2)=2(y+4)+2, (Apply distributive property)
4y-8=2y+8+2
-2y+8 -2y+8 ( Add -2y+8 to both sides and combine like terms)
--------------
2y=18 (Divide both sides by 2)

2y/2=18/2
y=9
Check:4(9-2)=2(9+4)+2
28=26+2
28=28
Done.